please help me solve this equiation:
sume methods
5x+4y=6
2x=-5y-1
thanks

Answer by checkley77(12844)

5x+4y=6
5x=-4y+6 multiply by 2
2x=-5y-1 multiply by -5 & add
10x=-8y+12
-10x=+25y+5
-----------------
0=17y+1717y=-17
y=-17/7
y=-1 ans.
5x+4*-1=6
5x-4=6
5x=6+4
5x=10
x=10/5
x=2 ans.
Proof:
2*2=-5*-1-1
4=5-1
4=4